# Charging a Quantum Battery Mediated by Parity-Deformed Fields

### v1

@@ 1 Introduction
In recent years, there has been a significant evolution in our understanding of energy harnessing, all thanks to the remarkable advancements achieved in the field of quantum thermodynamics. This has led to the emergence of a new concept called quantum batteries (QBs) [1], small-scale energy storage devices. QBs are quantum mechanical systems that can store energy temporarily in the finite dimensional degree of freedom using the principles of quantum mechanics [2, 3, 4, 5]. Compared to electrochemical batteries, QBs can potentially store much higher amounts of energy [6, 7, 8, 9, 10, 11] and experience incredibly fast charging times [12, 13, 14, 15, 16] because they take advantage of unique features like quantum coherence and quantum entanglement. An empty QB is usually charged based on an interaction protocol between QB itself with either an external field or a quantum system which serves as a charger. During the interaction, it moves from a lower energy level to a higher one and gets charged. Over the past years, many theoretical studies have been devoted on the different aspects of QBs. These studies include proposing charging processes [17, 18, 19, 20, 21, 22, 23, 24, 25], evaluating the performance of QBs analyzing some figures of merit such as the stored energy, ergotropy (the maximum energy extractable through cyclic unitary transformations) and charging efficiency [26, 27, 28, 29, 30]. Furthermore, different experimental methods have been suggested to generate QBs using the trapped ions [22, 31, 32], cold atoms [33], and superconducting qubits [34, 35, 36].
@@
Along with these developments, due to the inevitable interaction between a real quantum system and its environment, the destructive effects of environmental noises on the QB’s charging and discharging processes have been investigated [37, 38, 39]. Nowadays there is a growing fascination with the study QBs from the perspective of open quantum systems. When a QB is subjected into the environmental noises, its coherence leaks into the environment, resulting in a phenomenon known as decoherence. Unfortunately, the environment induced decoherence often reduces the performance of QBs during charging and discharging [40, 41]. In particular, decoherence pushes QBs towards a non-active (passive) equilibrium state during the charging process, making it extremely challenging to extract work from the QBs in a cyclic unitary process [42]. Decoherence can also affect QBs disconnected from both charger and consumption hub, leading to self-discharge of QBs [43, 44, 45]. Hence, designing QBs robust to the environmental dissipations is very important, as this is a vital step towards practical implementation of QBs in everyday life. Recent research efforts have focused not only on studying the effect of the environment on QBs decoherence effects [46, 47, 48, 49], but also developing open QBs that utilize quantum control techniques to enhance the stability of their charging performance [50, 51, 52]. For instance, in [51] the authors use dark states to achieve a high performance, robust QB comprising an ensemble of $1/2$-spins that are charged by another ensemble of $1/2$-spins inside a thermal reservoir. In another work, Centrone et al [52], proposed a continuous variable QB coupled weakly to a squeezed thermal reservoir and managed to control the steady-state charging performance of the battery by boosting the quantum squeezing of reservoir. A stabilization scheme based on a sequence of repeated quantum measurements has been proposed in [53]. Kamin et al [50] proposed a qubit-based QB that is charged by the mediation of a non-Markovian environment. They showed that the memory effects are beneficial for improving charging cycle performance. In addition to the above robust protocols, several other control techniques have been exploited to protect the charging cycle of QBs such as feedback control technique [54, 55, 56], Bang-Bang modulation of the intensity of an external Hamiltonian [57], convergent iterative algorithm [58], inhiring an auxiliary quantum system [59], modulating the detuning between system and reservoir [60], three level stimulated Raman adiabatic passage technique [61], engineering quantum environments [62, 48], etc.
@@
On the other hand, according to the previous studies on the Markovian and non-Markovian dynamics of open quantum systems, employing the technique of reservoir engineering through the utilization of leaky cavities with parity-deformed fields provides novel insights into control methods for stabilizing the entanglement and coherence of a two-qubit system against the environmental-induced dissipations [63, 64]. It has been shown that parity deformation of environment fields has significant capabilities in the controlling initial entanglement of the light-matter systems. The reason behind these capabilities is that parity deformation introduces field nonlinearities inside the environment and also induces an intensity-dependent coupling between the atom and environment that together control dynamics of the system [64]. In the last years, several works have been devoted to explore the role of parity deformation on the dynamics of light-matter systems [65, 66, 67, 69, 68]. In Ref [69], for example, it has been proven that parity-deformed fields can enable robust transfer of entangled states to far qubits of a cavity-based quantum network. It also has been shown that parity deformation can improve the power and efficiency of a quantum heat engine [68]. In this study, we would like to use the controlling capability of the parity-deformed fields to boost the charging performance of open qubit-based quantum batteries. Our attention is directed towards a wireless-like charging process, where the charger and battery’s qubits are immersed in a zero-temperature environment modeled by a collection of the parity-deformed field modes, and battery is charged by the mediation of the environment without needing an external energy source or direct connection to the charger. More precisely, we will investigate environment-mediated charging performance of the battery in both the Markovian (memoryless) and non-Markovian (memory-keeping) dynamical regimes, with emphasis on the control role of the induced field nonlinearity and the intensity-dependent qubit-environment coupling. We will show that quantum memory is crucial in improving charging performance of QB, and will prove that parity deformation is able to induce memory effects in the dynamics and to transform the Markovian process to the non-Markovian one.
The rest of the paper has been structured as follows. The physical model, open dynamics of the charger-battery system and an explicit expression of their evolved reduced density matrix are given in the second section. Section 3 is devoted to introduce and describe several figures of merit for characterizing the performance of QBs. Section 4 discuss our results. Finally, section 5 concludes this paper.
@@ 2 Model and Solution
Let us consider an open charger-battery system consisting of two qubits, the qubit A as a charger and the qubit B as a QB, both coupled to a zero-temperature reservoir (shown in Fig. 1). Each qubit is modeled by a two-level system with the ground state $|g\rangle$ and excitation state $|e\rangle$, but the reservoir is modeled by a collection of the parity-deformed field modes. In our system, there is no interaction between the qubits, and the battery is charged by the mediation of the reservoir, which under the rotating wave approximation is ruled by the Hamiltonian ($\hbar=1$):
@@
In the above set of equations, $\hat{\sigma}^{(i)}_{+}=|e\rangle_{i}\,{}_{i}\langle g|$, $\hat{\sigma}^{(i)}_{-}=|g\rangle_{i}\,{}_{i}\langle e|$ and $\hat{\sigma}^{(i)}_{z}=\left(|e\rangle_{i}\,{}_{i}\langle e|-|g\rangle_{i}\,{}_{i}\langle g|\right)$ $(i=A,B)$ are, respectively, arising, lowering and inversion operators of the qubit $i$ with the energy $\omega_{0}$. Moreover, $\hat{\mathbf{a}}_{k}$ and ${\hat{\mathbf{a}}^{{\dagger}}}_{k}$ are respectively the annihilation and creation operator of the $\mathit{k}$-mode of the reservoir para-Bose field with frequency $\omega_{k}$, which together with the parity operator $\hat{R}_{k}$ constitute a deformed algebra, namely, Wigner algebra [70, 71], with the following (anti-)commutation relations:
@@
where the real constant $\nu\in(-0.5,\infty)$ denotes the parity deformation parameter. The Wigner algebra can be represented by the generalized Fock states $\{\|n\rangle|_{n=0}^{\infty}\}$, as below
@@
Such a decomposition enables us to restrict dynamics of the whole qubit-reservoir system to the excitation subspaces. Here we are interested to restrict dynamics of the entire system up to the one-excitation subspace, i.e. $\mathcal{H}_{0}\oplus\mathcal{H}_{1}$. We choose a normalized initial state of whole qubit-reservoir system belong to the one-excitation subspace $\mathcal{H}_{1}$ as a linear superposition of vectors $|e_{A},g_{B}\rangle|0\rangle_{\mathcal{E}}$ and $|g_{A},e_{B}\rangle|0\rangle_{\mathcal{E}}$, as follows
@@
where $W$ is proportional to the vacuum Rabi frequency and $\Delta$ denotes the detuning of the atomic frequency $\omega_{0}$ and the central frequency of the cavity. In addition, $\lambda$ specifies the spectral width of the coupling which is linked to the memory time $\tau_{E}$ by the relation $\tau_{E}=\lambda^{-1}$. The relative magnitudes of $W$ and $\lambda$, is used to distinguish the weak coupling regime from the strong one [75]. For $\lambda>2W$, we have a weak coupling regime and dynamics of the charger-battery system is Markovian. In this regime information or energy decays irreversibly. However the strong coupling regime corresponds to $\lambda<2W$, where the dynamics is deemed non-Markovian. In the strong regime, the information or energy flows back from the reservoir to the system [73].
@@ 3 Energetics of the battery
where minimization in Eq. (22b) is taken over all the unitary transformations $U$ acting locally on the $\rho_{B}(t)=\mathrm{tr}\big{(}\rho_{AB}(t)\big{)}$. By considering the spectral decomposition of the Hamiltonian $H_{B}$ of the battery: $H_{B}|\varepsilon_{i}\rangle=\varepsilon_{i}|\varepsilon_{i}\rangle$ with the ordered eigenvalues as $\varepsilon_{i}\leq\varepsilon_{i+1}$, and the instantaneous spectral decomposition $r_{i+1}\leq r_{i}$ of the instantaneous battery state $\rho_{B}(t)$, associated to eigenvectors $|r_{i}(t)\rangle$, the ergotropy can be re-expressed as
@@ 4 Results and discussion
In this section, we will analyze the Markovian and non-Markovian charging performance of the introduced open QB. In particular, we investigate the role of parity deformation of mediator environment on the dynamical behavior of performance indicators including stored energy, ergotropy and efficiency. In what follows, we consider an initial condition in which the battery is empty and the charger has the maximum energy, i.e. $c_{1}(0)=1$, $c_{2}(0)=0$.
In Fig. 2, we plot the stored energy $\Delta E_{B}$ as a function of $\lambda t$ for different values of the deformation parameter $\nu$ by fixing $\Delta=\frac{2\nu}{2\nu+1}\omega_{0}$ and $r_{1}=r_{2}=1/\sqrt{2}$. In panel (a), the battery is charged by the mediation of a Markovian environment $(R=0.4)$, while in panel (b), it is charged by the mediation of a non-Markovian environment $(R=50)$. What is shown is that parity deformation of the mediator fields has positive impact in controlling the stored energy of battery in both the Markovian and non-Markovian charging regimes. In the Markovian scenario, as illustrated in panel (a), depending on whether $\nu=0$ or $\nu\neq 0$, different behaviors of $\Delta E_{B}$ emerge: in the $\nu=0$ case, $\Delta E_{B}$ shows a generally monotonic increasing behavior, and $\nu<0$ shows a non-monotonous behavior, similar to what can be seen in panel (b). The non-monotonic behavior as a manifestation of non-Markovian dynamics indicates that the parity deformation introduces an enhancement of memory effects, and causes a crossover from a Markovian to a non-Markovian in the dynamics. Mathematically, the origin of this crossover should be traced back to the survival amplitude $p(t)$ given in (19) because this function, which generally determines all possible trajectories of the system dynamics in the damped fashion, depends on the deformation parameter $\nu$ by the Eq. (20), and enhancement of memory effects due to the parity deformation of the mediator fields is potentially possible. We will investigate this possibility below in the strong coupling regime.
@@
In Fig. (3) we plot the $\mathcal{N}$ as function of deformation parameter $\nu$ for a pair of initial states $\rho_{1}(0)=|e_{A},g_{B}\rangle$, $\rho_{2}(0)=\frac{1}{\sqrt{2}}[|e_{A},g_{B}\rangle+|g_{A},e_{B}\rangle]$associated to the charger qubit $A$ and battery qubit $B$. Here we choose $R=0.4$ to ensure the memory effects of the mediator are not prominent. As shown in the figure, the system exhibits Markovian dynamics until the negative values of $\nu$ are small and below a certain threshold (e.g., $\nu=-0.3$ in the figure). However, when $\nu$ exceeds this threshold the Markovian dynamics of the charger-battery system changes to non-Markovian (i.e., $\mathcal{N}>0$) signifying enhancement of memory effects of the environmental noises. In this situation, non-Markovianity increases asymptotically by regular increasing negative values of the parity deformation parameter. Therefore, parity deformation is able to enhance the memory effects of the mediator environment on the charging process.
@@
Finally, we examine the effect of parity deformation of mediator fields on the dynamics of the battery efficiency. The results for Markovian and non-Markovian charging processes are illustrated in Fig. 5(a) and 5(b), respectively. Here we consider the same parameter values as in Fig. 2. Comparing Fig. 5 with 4, we can see that both efficiency and ergotropy are positively affected by the parity deformation of mediator fields. However the efficiency is influenced more than the ergotropy; the amount of increment in efficiency is more than the ergotropy in both Markovian and non-Markovian charging processes.

### v2

@@ 1 Introduction
In recent years, there has been a significant evolution in our understanding of energy harnessing, all thanks to the remarkable advancements achieved in the field of quantum thermodynamics. This has led to the emergence of a new concept called quantum batteries (QBs) [1], small-scale energy storage devices. QBs are quantum mechanical systems that can store energy temporarily in the finite dimensional degree of freedom using the principles of quantum mechanics [2, 3, 4, 5]. Compared to electrochemical batteries, QBs can experience fast charging times [6, 7, 8, 9, 10] because they take advantage of unique features like quantum coherence and quantum entanglement. An empty QB is usually charged based on an interaction protocol between QB itself with either an external field or a quantum system which serves as a charger. During the interaction, it moves from a lower energy level to a higher one and gets charged. The charging performance of a QB is characterized by its storage energy, efficiency, power, storage capacity as well as ergotropy (the maximum energy extractable from the battery through cyclic unitary). Quantum coherence and quantum entanglement have been proven to be beneficial in charging performance of QBs [11, 12, 13]. Furthermore, a powerful charging can be generated by taking into account the collective quantum effects of a battery with $N$ quantum cells [14, 15, 16, 17]. For such QBs, the total power scales with $N\sqrt{N}$, instead of (linear) $N$ [14]. Over the past years, many theoretical studies have been devoted on the different aspects of QBs. These studies include proposing charging processes [18, 19, 20, 21, 22, 23, 24, 25, 26, 27], evaluating the charging performance [28, 12, 29, 30, 31, 32], suggesting possible implementation schemes [24, 33, 34, 35, 36, 37, 38, 39, 40, 41], and so on. A comprehensive review on quantum batteries can be found in Ref. [42].
@@
Along with these developments, due to the inevitable interaction between a real quantum system and its environment, the destructive effects of environmental noises on the QB’s charging and discharging processes have been investigated [43, 44, 45]. Nowadays there is a growing fascination with the study QBs from the perspective of open quantum systems. When a QB is subjected into the environmental noises, its coherence leaks into the environment, resulting in a phenomenon known as decoherence. Unfortunately, the environment induced decoherence often reduces the performance of QBs during charging and discharging [46, 47]. In particular, decoherence pushes QBs towards a non-active (passive) equilibrium state during the charging process, making it extremely challenging to extract work from the QBs in a cyclic unitary process [48]. Decoherence can also affect QBs disconnected from both charger and consumption hub, leading to self-discharge of QBs [49, 50, 51]. Hence, designing QBs robust to the environmental dissipations is a key task. Recent research efforts have focused not only on studying the effect of the environment on QBs decoherence effects [52, 53, 54, 55], but also developing open QBs that utilize quantum control techniques to enhance the stability of their charging performance [56, 57, 58]. For instance, in [57] the authors use dark states to achieve a high performance, robust QB comprising an ensemble of $1/2$-spins that are charged by another ensemble of $1/2$-spins inside a thermal reservoir. In another work, Centrone et al [58], proposed a continuous variable QB coupled weakly to a squeezed thermal reservoir and managed to control the steady-state charging performance of the battery by boosting the quantum squeezing of reservoir. In [41] Dou and Yang proposed a superconducting transmon qubit-resonator QB consisting of $N$ coupled transmon qubits as the battery and a one-dimensional transmission line resonator as the charger. They showed that, when qubits undergo the dephasing channel, transitions between the qubit states occur and the decay of qubit energy into the environment is suppressed, so that a high performance QB is achieved. They also imposed the condition that the anharmonicity of transmon is smaller than the transmon-resonator detuning [59], and found that by tuning transmons into a decoherence-free subspace, a resonator-qutrits QB with a stable and efficient charging process can be realized. A stabilization scheme based on a sequence of repeated quantum measurements has been proposed in [60]. Kamin et al [56] proposed a qubit-based QB that is charged by the mediation of a non-Markovian environment. They showed that the memory effects are beneficial for improving charging cycle performance. In addition to the above robust protocols, several other control techniques have been exploited to protect the charging cycle of QBs such as feedback control technique [61, 62, 63], Bang-Bang modulation of the intensity of an external Hamiltonian [64], the inversely-engineered control based on Lewis-Riesenfeld invariants [65], convergent iterative algorithm [66], inhiring an auxiliary quantum system [67], modulating the detuning between system and reservoir [68], three level stimulated Raman adiabatic passage technique [69], engineering quantum environments [70, 54], etc.
@@
On the other hand, according to the previous studies on the Markovian and non-Markovian dynamics of open quantum systems, employing the technique of reservoir engineering through the utilization of leaky cavities with parity-deformed fields provides novel insights into control methods for stabilizing the entanglement and coherence of a two-qubit system against the environmental-induced dissipations [71, 72]. It has been shown that parity deformation of environment fields has significant capabilities in the controlling initial entanglement of the light-matter systems. The reason behind these capabilities is that parity deformation introduces field nonlinearities inside the environment and also induces an intensity-dependent coupling between the atom and environment that together control dynamics of the system [72]. In the last years, several works have been devoted to explore the role of parity deformation on the dynamics of light-matter systems [73, 74, 75, 76, 77]. In Ref [76], for example, it has been proven that parity-deformed fields can enable robust transfer of entangled states to far qubits of a cavity-based quantum network. It also has been shown that parity deformation can improve the power and efficiency of a quantum heat engine [77]. At this point it is worth to remark that the para-particles, despite being theoretically well defined, have not yet been experimentally confirmed. Recently, some effective methods for simulating para-particles have been proposed in the literature [78, 79, 80, 81]. Particularly, a quantum system consisting of a qubit with two field modes in strong coupling regime has been considered as a promising physical system for simulating the parity-deformed fields [80]. Also, one-dimensional arrays of coupled waveguides [81] is another interesting physical system that could offer a classical simulation of the parity-deformed fields by taking advantage of their f-deformed realization.
In this study, we would like to use the controlling capability of the parity-deformed fields to boost the charging performance of open qubit-based quantum batteries. Our attention is directed towards a wireless-like charging process, where the charger and battery’s qubits are immersed in a zero-temperature environment modeled by a collection of the parity-deformed field modes, and battery is charged by the mediation of the environment without needing an external energy source or direct connection to the charger. More precisely, we will investigate environment-mediated charging performance of the battery in both the Markovian (memoryless) and non-Markovian (memory-keeping) dynamical regimes, with emphasis on the control role of the induced field nonlinearity and the intensity-dependent qubit-environment coupling. We will show that quantum memory is crucial in improving charging performance of QB, and will prove that parity deformation is able to induce memory effects in the dynamics and to transform the Markovian process to the non-Markovian one. The rest of the paper has been structured as follows. The physical model, open dynamics of the charger-battery system and an explicit expression of their evolved reduced density matrix are given in the second section. Section 3 is devoted to introduce and describe several figures of merit for characterizing the performance of QBs. Section 4 discuss our results. Finally, section 5 concludes this paper.
@@ 2 Model and Solution
Let us consider an open charger-battery system consisting of two qubits, the qubit A as a charger and the qubit B as a QB, both coupled to a zero-temperature reservoir (shown in Fig. 1). Each qubit is modeled by a two-level system with the ground state $|g\rangle$ and excited state $|e\rangle$, but the reservoir is modeled by a collection of the parity-deformed field modes. In our system, there is no interaction between the qubits, and the battery is charged by the mediation of the reservoir, which under the rotating wave approximation is ruled by the Hamiltonian ($\hbar=1$):
@@
In the above set of equations, $\hat{\sigma}^{(i)}_{+}=|e\rangle_{i}\,{}_{i}\langle g|$, $\hat{\sigma}^{(i)}_{-}=|g\rangle_{i}\,{}_{i}\langle e|$ and $\hat{\sigma}^{(i)}_{z}=\left(|e\rangle_{i}\,{}_{i}\langle e|-|g\rangle_{i}\,{}_{i}\langle g|\right)$ $(i=A,B)$ are, respectively, arising, lowering and inversion operators of the qubit $i$ with the energy $\omega_{0}$. $\hat{\mathbf{a}}_{k}$ and ${\hat{\mathbf{a}}^{{\dagger}}}_{k}$ are respectively the annihilation and creation operator of the $\mathit{k}$-mode of the reservoir para-Bose field with frequency $\omega_{k}$. Moreover, $\alpha_{i}|g_{k}|$ denotes the coupling strength of the qubit $i$ with mode $k$ of the parity deformed field. The the annihilation and creation operator $\hat{\mathbf{a}}_{k}$ and ${\hat{\mathbf{a}}^{{\dagger}}}_{k}$ together with the parity operator $\hat{R}_{k}$ constitute a deformed algebra, namely, Wigner algebra [82, 83], with the following (anti-)commutation relations:
@@
where the real constant $\nu\in(-0.5,\infty)$ denotes the parity deformation parameter. Note that, for $\nu=0$, the Wigner algebra is reduced to the Weyl-Heisenberg algebra of the bosonic filed modes. The Wigner algebra can be represented by the generalized Fock states $\{\|n\rangle|_{n=0}^{\infty}\}$, as below
@@
Such a decomposition enables us to restrict dynamics of the whole qubit-reservoir system to the excitation subspaces. Here we are interested to restrict dynamics of the entire system to the one-excitation subspace, i.e. $\mathcal{H}_{1}$. We choose a normalized initial state of whole qubit-reservoir system belong to the one-excitation subspace $\mathcal{H}_{1}$ as a linear superposition of vectors $|e_{A},g_{B}\rangle|0\rangle_{\mathcal{E}}$ and $|g_{A},e_{B}\rangle|0\rangle_{\mathcal{E}}$, as follows
@@
The restriction of dynamics to the single-excitation subspace enable us to simulate the Hamiltonian (5) as a spin model consisting of two uncoupled spins surrounded by a chain of other spins. To show this, inspired by the idea given in [85, 86], we identify the qubit and reservoir operators in Hamiltonian (5) as Pauli operators, and encode the subspace $\mathcal{H}_{1}$ to the Hilbert space of the spin model. The qubit part is trivially mapped to spins. For the reservoir, we use a direct one-to-one mapping and encode the bosonic basis $|0\rangle_{\mathcal{E}}$ and $|1_{k}\rangle_{\mathcal{E}}$, as follows
$|0\rangle_{\mathcal{E}}=|0,0,...\rangle_{\mathcal{E}}\rightarrow|\downarrow\rangle_{1}\otimes|\downarrow\rangle_{2}\otimes...,$ (10a)
$|1_{k}\rangle_{\mathcal{E}}=|0,...,0_{k-1},1_{k},0_{k+1},0,..\rangle_{\mathcal{E}}\rightarrow|\downarrow\rangle_{1}\otimes...\otimes|\downarrow\rangle_{k-1}\otimes|\uparrow\rangle_{k}\otimes|\downarrow\rangle_{k+1}\otimes...,$ (10b)
where $|\downarrow\rangle_{k}=\left(\begin{array}[]{c}0\\ 1\\ \end{array}\right)$ and $|\uparrow\rangle_{k}=\left(\begin{array}[]{c}1\\ 0\\ \end{array}\right)$ are spinors of the k-th spin of the chain. Furthermore, we use the actions of the field’s operators on the subspace $\mathcal{H}_{1}$ and map them to generators of the Pauli group,
$a_{k}\rightarrow|\downarrow\rangle_{k}\,{}_{k}\langle\uparrow|\equiv\hat{\sigma}^{(k)}_{-},$ (11a)
$(-1)^{n_{k}}\rightarrow|\uparrow\rangle_{k}\,{}_{k}\langle\uparrow|-\downarrow\rangle_{k}\,{}_{k}\langle\downarrow|\equiv\hat{\sigma}^{(k)}_{z}.$ (11b)
Now, according to Eqs. (11a) and (11b), the Hamiltonian (5) can be rewritten as Hamiltonian of open quantum system consisting of two central spin $A$ and $B$ that are surrounded by a spin environment,
$\frac{\omega_{0}}{2}\sum_{i=A,B}\hat{\sigma}^{(i)}_{z}+(2\nu+1)\sum_{k}\omega_{k}\hat{\sigma}^{(k)}_{+}\hat{\sigma}^{(k)}_{-}+\sqrt{2\nu+1}\sum_{i=A,B}\sum_{k}\big{(}\alpha_{i}g_{k}\hat{\sigma}^{(k)}_{-}\hat{\sigma}^{(i)}_{+}+H.c.\big{)}.$ (12)
Here the field nonlinearity $\big{(}1-(-1)^{\hat{n}_{k}}\big{)}$ and intensity-dependent coupling $\big{(}\hat{a}_{k}F(\hat{n}_{k})\hat{\sigma}^{(i)}_{+}+H.c.\big{)}$ have been identified respectively as $\hat{\sigma}^{(k)}_{+}\hat{\sigma}^{(k)}_{-}$ and $\big{(}\hat{\sigma}^{(k)}_{-}\hat{\sigma}^{(i)}_{+}+H.c.\big{)}$ interactions which can be potentially implemented by the superconducting integrated circuits [87, 88].
@@
where $W$ is proportional to the vacuum Rabi frequency and $\Delta$ denotes the detuning of the atomic frequency $\omega_{0}$ and the central frequency of the reservoir. In addition, $\lambda$ specifies the spectral width of the coupling which is linked to the memory time $\tau_{E}$ by the relation $\tau_{E}=\lambda^{-1}$. The relative magnitudes of $W$ and $\lambda$, is used to distinguish the weak coupling regime from the strong one [91]. For $\lambda>2W$, we have a weak coupling regime and dynamics of the charger-battery system is Markovian. In this regime information or energy decays irreversibly. However the strong coupling regime corresponds to $\lambda<2W$, where the dynamics is deemed non-Markovian. In the strong regime, the information or energy flows back from the reservoir to the system [89].
@@ 3 Energetics of the battery
where minimization in Eq. (25b) is taken over all the unitary transformations $U$ acting locally on the $\rho_{B}(t)=\mathrm{tr}_{A}\big{(}\rho_{AB}(t)\big{)}$. By considering the spectral decomposition of the Hamiltonian $H_{B}$ of the battery: $H_{B}|\varepsilon_{i}\rangle=\varepsilon_{i}|\varepsilon_{i}\rangle$ with the ordered eigenvalues as $\varepsilon_{i}\leq\varepsilon_{i+1}$, and the instantaneous spectral decomposition $r_{i+1}\leq r_{i}$ of the instantaneous battery state $\rho_{B}(t)$, associated to eigenvectors $|r_{i}(t)\rangle$, the ergotropy can be re-expressed as
@@ 4 Results and discussion
In this section, we will analyze the Markovian and non-Markovian charging performance of the introduced open QB. In particular, we investigate the role of parity deformation of mediator environment as well as the coupling strength of qubits with the mediator on the dynamical behavior of performance indicators including stored energy, ergotropy and efficiency. In what follows, we consider an initial condition in which the battery is empty and the charger has the maximum energy, i.e. $c_{1}(0)=1$, $c_{2}(0)=0$.
@@
In Fig. 2, we plot the stored energy $\Delta E_{B}$ as a function of the dimensionless quantity $\lambda t$ for different values of the deformation parameter $\nu$ by choosing $\Delta=\frac{2\nu}{2\nu+1}\omega_{0}$ and $r_{1}=r_{2}=1/\sqrt{2}$. The rationale for choosing $r_{1}=r_{2}=1/\sqrt{2}$ is that, as we will show later, the QB’s performance is optimized only for these specific parameter values. On the other hand, from Eqs. (22) and (23) one can check that by setting $\Delta=\frac{2\nu}{2\nu+1}\omega_{0}$ charging process of the QB is acutely independent of the qubit’s transition frequency $\omega_{0}$ and thus our following results will be valid in different regimes of parameter. Furthermore, since in Ref. [56] the charger and battery’s qubits are symmetrically coupled to the environment ($r_{1}=r_{2}=1/\sqrt{2}$) and also the qubit’s transition frequency is in resonant with the central frequency of environmental modes ($\Delta=0$), its results can be now recovered in the limit $\nu=0$. In panel (a), the battery is charged by the mediation of a Markovian environment $(R=0.4)$, while in panel (b), it is charged by the mediation of a non-Markovian environment $(R=50)$. What is shown is that parity deformation of the mediator fields has positive impact in controlling the stored energy of battery in both the Markovian and non-Markovian charging regimes. In the Markovian scenario, as illustrated in panel (a), depending on whether $\nu=0$ or $\nu\neq 0$, different behaviors of $\Delta E_{B}$ emerge: in the $\nu=0$ case, $\Delta E_{B}$ shows a generally monotonic increasing behavior, and $\nu<0$ shows a non-monotonous behavior, similar to what can be seen in panel (b). The non-monotonic behavior as a manifestation of non-Markovian dynamics indicates that the parity deformation introduces an enhancement of memory effects, and causes a crossover from a Markovian to a non-Markovian in the dynamics. Mathematically, the origin of this crossover should be traced back to the survival amplitude $p(t)$ given in (22) because this function, which generally determines all possible trajectories of the system dynamics in the damped fashion, depends on the deformation parameter $\nu$ by the Eq. (23), and enhancement of memory effects due to the parity deformation of the mediator fields is potentially possible. We will investigate this possibility below in the strong coupling regime.
Figure 4: Dynamics of the stored energy $\Delta E_{B}(t)$ as function of the relative coupling strength $r_{1}$ for (a) Markovian dynamics with $R=0.4$ and (b) non-Markovian dynamics with $R=50$ by setting $\nu=-0.3$. The other parameters are taken as in Fig. 2.
@@
In Fig. (3) we plot the $\mathcal{N}$ as function of deformation parameter $\nu$ for an optimal pair of initial state associated to the charger-battery system. Here we choose $R=0.4$ to ensure the memory effects of the mediator are not prominent. As shown in the figure, the system exhibits Markovian dynamics until the negative values of $\nu$ are small and below a certain threshold (e.g., $\nu=-0.3$ in the figure). However, when $\nu$ exceeds this threshold the Markovian dynamics of the charger-battery system changes to non-Markovian (i.e., $\mathcal{N}>0$) signifying enhancement of memory effects of the environmental noises. In this situation, non-Markovianity increases asymptotically by regular increasing negative values of the parity deformation parameter. Therefore, parity deformation is able to enhance the memory effects of the mediator environment on the charging process.
@@
Now, we examine the influence of the relative coupling strength $r_{1}$ on the charging performance of QB. In Fig. 4 we plot the stored energy $\Delta E_{B}$ as a function of $\lambda t$ and $r_{1}$ in the Markovian (panel (a)) and non-Markovian (panel (b)) charging processes. This figure clearly shows that, for $r_{1}=0$ and 1, no energy stored in the QB as these correspond to cases where only one qubit interacts with the mediator environment and there is no correlation between qubits due to the environment. In both the Markovian and non-Markovian charging processes, the maximum value of stored energy is achieved at $r_{1}\approx 0.71=\frac{1}{\sqrt{2}}$, i.e., when the charger and battery’s qubits are symmetrically coupled to the environment. Comparing panel (a) with (b) reveals fundamental differences of the Markovian and non-Markovian charging processes. In the Markovian charging process, $\Delta E_{B}$ always increases monotonically in time and reaches a steady-state value, while in the non-Markovian process it experiences an oscillatory behavior due to the memory effects of the environment. We find that the memory effects improve energy storage and significantly increase the charging speed. In the Markovian charging process, the battery can be charged up to a $30\%$ of its capacity, and the optimal time for this charge is about $\lambda t\approx 15$. However, in the non-Markovian process the battery can be full charged, and optimal charging time is about $\lambda t\approx 0.1$.
@@
Finally, we examine the effect of parity deformation of mediator fields on the dynamics of the battery efficiency. The results for Markovian and non-Markovian charging processes are illustrated in Fig. 6(a) and 6(b), respectively. Here we consider the same parameter values as in Fig. 2. Comparing Fig. 6 with 5, we can see that both efficiency and ergotropy are positively affected by the parity deformation of mediator fields. However the efficiency is influenced more than the ergotropy; the amount of increment in efficiency is more than the ergotropy in both Markovian and non-Markovian charging processes.